# Asking Fast Radio Bursts (FRBs) for More than Reionization History

### v1

@@ Abstract
We propose different estimators to probe the epoch of reionization (EoR) intergalactic medium (IGM) using the dispersion measure (${\rm DM}$) of the FRBs. We consider three different reionization histories which we can distinguish with a total of $\lesssim 1000$ ${\rm DM}$ measurements during EoR if their redshifts are known. We note that the redshift derivatives of ${\rm DM}$ are also directly sensitive to the reionization history. The major point of this work is exploring the variance in the ${\rm DM}$ measurements and the information encoded in them. We find that the all-sky average $\overline{{\rm DM}}(z)$ gets biased from the LoS fluctuations in the ${\rm DM}$ measurements introduced by the ionization of IGM during EoR. We find that the ratio $\sigma_{\rm DM}/\overline{{\rm DM}}$ depends directly on the ionization bubble sizes as well as the reionization history. On the other hand, we also find that angular variance (coined as structure function) of ${\rm DM}$ encodes the information about the duration of reionization and the typical bubble sizes as well. We establish the usefulness of variances in ${\rm DM}$ using toy models of reionization and later verify it with the realistic reionization simulations.
@@ 1 Introduction
According to the current understanding of cosmology, our universe transitioned from being a highly cold-neutral phase in the past to almost hot-ionized phase at present. This is supposed to be a result of UV radiations from the very first objects that formed in the universe photoionizing the intergalactic medium (IGM) (see e.g., Loeb & Barkana, 2001; Zaroubi, 2013; Pritchard & Loeb, 2012). This window of transition is termed as the Epoch of Reionization (EoR). The study of the EoR is crucial for answering several questions regarding the emergence of the first sources, their properties and impact on the IGM, evolution to present day structures, the exact timeline of this epoch, etc. Despite our efforts in the last few decades, our understanding of the EoR remains limited (see e.g., Shaw et al., 2023).
Our present understanding of the timing and duration of EoR is guided by a few indirect observations such as the measurements of the Thomson scattering optical depth from the cosmic microwave background radiation observations (e.g., Planck Collaboration et al., 2020) and the Gunn-Peterson troughs in the high-$z$ quasar spectra (e.g., Becker et al., 2001; Fan et al., 2006; Gallerani et al., 2006; Becker et al., 2015; Bosman et al., 2022; D’Odorico et al., 2023; Gaikwad et al., 2023; Greig et al., 2024; Asthana et al., 2024; Spina et al., 2024). Additional constraints on the timeline of EoR come from the recent observations of the high-$z$ Ly-$\alpha$ emitters (e.g., Hu et al., 2010; Kashikawa et al., 2011; Ota et al., 2017; Ishigaki et al., 2018; Morales et al., 2021; Bruton et al., 2023; Nakane et al., 2023) and their clustering measurements (e.g., Faisst et al., 2014; Santos et al., 2016; Wold et al., 2022), Lyman break galaxies (Mason et al., 2018; Hoag et al., 2019; Naidu et al., 2020; Bolan et al., 2022), and the Ly-$\alpha$ damping wings in the high-$z$ quasar spectra (e.g., Bañados et al., 2018; Davies et al., 2018; Ďurovčíková et al., 2020; Wang et al., 2020; Yang et al., 2020; Umeda et al., 2023; Ďurovčíková et al., 2024). These experiments attempt to constrain the reionization history by putting bounds on the global ionization fraction of the IGM during EoR. On the other hand, the measurements of the effective optical depth of Ly-$\alpha$ forests (using dark-gap/pixel statistics) (e.g., McGreer et al., 2014; Keating et al., 2019; Kulkarni et al., 2019; Zhu et al., 2021, 2022; Bosman et al., 2022) suggests that the end of reionization has a longer tail extending to somewhere between $z=5.5$ and $5.0$ instead of $z\approx 6$. However, all these analyses are either model-dependent or suffer from statistical variance, thus providing only loose bounds on the EoR timeline.
@@
Probing EoR directly using the redshifted 21-cm signal with the current instruments is also challenging because of several hindrances such as large ($\sim 10^{4}$ times) foregrounds (e.g., Ali et al., 2008; Bernardi, G. et al., 2009, 2010; Ghosh et al., 2012), thermal noise, radio frequency interference, ionospheric turbulence and other systematics. While no undisputed detection of the EoR 21-cm signal has been achieved so far, the current data from the radio-interferometric experiments have been able to provide a few upper limits on the EoR 21-cm power spectra (e.g., LOFAR: Patil et al. 2017; Mertens et al. 2020, MWA: Barry et al. 2019; Li et al. 2019; Trott et al. 2020, HERA: Abdurashidova et al. 2022; The HERA Collaboration et al. 2022), and the upper-limits are improving gradually.
@@
A few earlier works have demonstrated the potential of highly energetic astrophysical events such as gamma-ray bursts (GRBs) (e.g., Ciardi & Loeb, 2000; Ioka, 2003; Inoue, 2004; Lidz et al., 2021) and fast radio bursts (FRBs) (e.g., Beniamini et al., 2021; Hashimoto et al., 2021; Heimersheim et al., 2022) during EoR as a probe to measure the reionization history. In this work we focus on the FRBs, which are luminous short-duration ($\sim$ few ms) astrophysical radio pulses which have been detected within a frequency band of $0.1-7~{}{\rm GHz}$ (see Petroff et al., 2022, for a review). Empirical studies on the all-sky event rates of FRBs based on observations find it to be $\sim 10^{3}/$day above the fluence limit of $5$ Jy ms and at a central frequency of $500$ MHz (The CHIME/FRB Collaboration et al., 2021), and the rate is expected to increase significantly at lower fluence thresholds.
@@
FRB signals disperse while travelling through the ionized medium. The amount of dispersion, quantified as Dispersion Measure (DM), directly depends on the free electron content along its path. The DM of a cosmological FRB is expected to have dominant contribution coming from the electrons in the IGM. During post-EoR ($z\lesssim 5.5$), where the IGM is almost ionized, the IGM DM roughly scales directly with the distances. Therefore the DM measurements can be turned to infer the redshift-distance of the FRB (Zhang, 2018; Kumar & Linder, 2019). Conversely, knowing the redshift of the FRBs accurately can be potentially used to estimate baryonic content of IGM during the post-EoR (e.g., McQuinn, 2013; Macquart et al., 2020; Lee et al., 2022; Khrykin et al., 2024), probe the epoch of second helium reionization (e.g., Caleb et al., 2019; Linder, 2020; Bhattacharya et al., 2021; Lau et al., 2021) and constrain several cosmological parameters (e.g., Deng & Zhang, 2014; Zhou et al., 2014; Yang & Zhang, 2017; Walters et al., 2018; Pol et al., 2019; Jaroszynski, 2019; Wu et al., 2020; Wucknitz et al., 2021; James et al., 2022; Hagstotz et al., 2022). In this work we explore how useful they can be as detailed probes of the EoR.
@@
Despite their enigmatic origin, a recent discovery of a galactic FRB (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020; Bera et al., 2024) clearly associates at least some FRBs to magnetars. Other, less direct, evidence linking FRBs to magnetars comes from the statistical properties of the bursts, from host galaxies and offsets relative to them and from the energetics and temporal properties of the bursts (e.g., Popov & Postnov, 2010; Falcke, Heino & Rezzolla, Luciano, 2014; Ioka, 2020; Wadiasingh et al., 2020; Zhang, 2022; Beniamini & Kumar, 2023; Chen et al., 2023; Totani & Tsuzuki, 2023). Hence, we can expect a sufficiently large number of FRBs during EoR ($z>6$) which spans a much larger time compared to the life-time of the massive Pop III stars ($\sim~{}5-30~{}{\rm Myr}$) that leave behind neutron star (NS) remnants with large angular momentum and strong magnetic fields. There are indirect evidences which supports a large abundance of high-$z$ FRBs (see the introduction of Beniamini et al., 2021, for more details).
@@
With the possibility of detecting high-$z$ FRBs (e.g., Hashimoto et al., 2020), one can turn their precise DM measurements to probe the sources and IGM during reionization. Recently, a few theoretical studies (Beniamini et al., 2021; Hashimoto et al., 2021; Heimersheim et al., 2022) have demonstrated the feasibility of using the DM measurements to extract the reionization history and Thomson scattering optical depth ($\tau_{\rm Th}$). Pagano & Fronenberg (2021) and Maity (2024) have used the mean DM from their synthetic FRB population to constrain the parameters of their reionization simulation. The results in most of these works are based on the assumption of knowing the precise redshift (spectroscopic or empirically) of FRBs. However, detecting the precise spectroscopic redshifts from the host galaxies is a challenging task with the current instruments. Conversely, Beniamini et al. (2021) suggest that the maximum value of DM for bursts spanning the EoR can provide an independent estimate of the Thomson optical depth of the universe without requiring direct redshift information. They have also shown that $\sim 40$ FRBs are sufficient to estimate average electron fraction in $4$ $z$-bins (within $z=6$ to $10$) with $4\%$ accuracy, if their redshifts could be determined with $10\%$ uncertainty. Similar results have been reported in Heimersheim et al. (2022) where they also estimated $\tau_{\rm Th}$ and the mid-redshift of reionization using the average DM. Beniamini et al. (2021) also suggested that the reionization history can be constrained from the determination of the number of FRBs during the EoR per unit DM, i.e., $dN_{\rm FRB}/d{\rm DM}$. Whereas, Hashimoto et al. (2021) have shown that the redshift derivatives of DM have potential to directly constrain the reionization history.
@@
In this work, we forward model the FRB signal measurements using estimators like globally-averaged DM, its redshift derivative, global dispersion and angular dispersion (along different LoS). For the first time, we demonstrate that fluctuations in the LoS in DM significantly impact measurements during the EoR, potentially biasing mean DM estimates compared to those derived using the mean ionization fraction of the IGM. We also demonstrate that $d{\rm DM}/dz$ and the scatter in DM along different LoS (as a function of $z$) has the potential to discern different reionization histories and morphologies of the IGM. We explore a novel aspect of angular dispersion in DM (defined as structure function in §2.2) at different redshifts. This structure function encodes information regarding the typical bubble sizes in the IGM as it probes the angular fluctuations. We validate our estimators on simple toy models of EoR light-cone signals. Later, we also apply our estimators to more realistic light-cones obtained from simulations. We also perform a comparative study between different reionization histories. For this study, we primarily assume a scenario where the redshifts of the FRBs are known to within $10\%$ uncertainty. Later, we ignore the redshift information and compute the marginalized average DM over the EoR window.
@@
This manuscript is arranged as follows. We define DM and structure function estimators in §2. In §3, we validate our estimators with the toy ionization field model. This is followed by §4 where we briefly describe the details of the actual EoR simulations, and the present the corresponding results under its subsection where we discuss the impact of post-EoR IGM on the DM estimates and presented results. Finally, we summarize and conclude this exercise in §5. The cosmological simulation here uses the cosmological parameter values $\Omega_{\rm m}=0.27$, $\Omega_{\Lambda}=0.73$, $\Omega_{\rm b}=0.044$ and $h=0.7$ adapted from Hinshaw et al. (2013).
@@ 2 Methodology
We revisit the basic theory of the IGM DM in the context of FRBs and define DM estimators which will be used in this work.
@@ 2.1 Dispersion Measure
The multifrequency FRB pulses disperse while traveling through an ionized medium due to their interaction with the free electrons along the way. The time delay in the arrival of the signal at frequency $\nu$, $\Delta t\propto\nu^{-2}{\rm DM}$; where DM is the line-integral of the free electron density. The total observed time-delay/DM will have contributions from the host galaxy ($\rm{DM}_{host}$), the Milky Way galaxy including the halo ($\rm{DM}_{MW}$), and the IGM (${\rm{DM}_{IGM}}$). For $\rm{DM}_{MW}$ we have reasonably good Galactic maps, and this can be largely removed from the data. Further, $\rm{DM}_{host}$ is reduced by a factor of $(1+z)^{-1}$ in the observer frame and so is suppressed when considering high-$z$ events, whereas ${\rm DM}_{\rm IGM}$ increases with $z$. Hence, in this work we only focus on studying ${\rm{DM}_{IGM}}$ and ignore any further discussion on other DM components, unless stated otherwise.
@@
[…]
where $c$ is speed of light in vacuum and $n_{e}(\bm{\theta},z)$ is the proper number density of free electrons. $H(z)=H_{0}\sqrt{\Omega_{m0}(1+z)^{3}+\Omega_{\Lambda 0}}$ denotes the Hubble parameter with $H_{0}$, $\Omega_{m0}$ and $\Omega_{\Lambda 0}$ respectively being the present day values of Hubble constant, matter density parameter and dark energy parameter. Baryons being the primary source of free electrons during and after EoR, we can write $n_{e}$ in terms of the baryon density parameter $\Omega_{\rm b0}$ and recast eq. (1) as
[…]
where $G$ is the gravitational constant, $m_{\rm H}$ is the mass of hydrogen atom, $\Delta$ is the matter overdensity and $x_{\rm{H}\,\textsc{ii}}$ denotes the ionization fraction of the IGM. Here, $\Delta$ includes the effects of evolution of the underlying matter density in the IGM whereas $x_{\rm{H}\,\textsc{ii}}$ is controlled by the photon field responsible for ionizing the IGM. On large scales, $x_{\rm{H}\,\textsc{ii}}=0$ before the reionization starts and it eventually approaches unity towards the end of EoR when the IGM is almost completely ionized. We obtain both $\Delta$ and $x_{\rm{H}\,\textsc{ii}}$ from our simulations which we discuss in a later section. The treatment of eq. (2) assumes the hydrogen and helium constitute almost the entire baryons in the universe, and the helium being $25\%$ of it by mass. Our model of IGM also assumes that the ionization of He i to He ii occurs concurrently with H i reionization.
The LoS path which the FRB signal traverses during post-reionization is effectively very large ($\sim 6000~{}{\rm Mpc}$), adding a larger contribution to total ${\rm{DM}_{IGM}}$. Post-EoR contribution acts as a nuisance since we are only interested in the impact of the reionization to the DM measurements. Hence, we restrict the lower limit of the integral (eq. 2) to the end of reionization $z_{\rm end}$ and define ${\rm{DM}_{EoR}}$:
[…]

### v2

# Asking Fast Radio Bursts for More than Reionization History

@@ Abstract
We propose different estimators to probe the intergalactic medium (IGM) during epoch of reionization (EoR) using the dispersion measure ($\rm DM$}) of the fast radio bursts. We consider three different reionization histories, which we can distinguish with a total of $\lesssim 1000\,{\rm DM}$ measurements during EoR if their redshifts are known. We note that the redshift derivatives of $\rm DM$} are also directly sensitive to the reionization history. The major point of this work is to explore the variance in the $\rm DM$} measurements and the information encoded in them. We find that the all-sky average $\overline{{\rm DM}}(z)$ gets biased from the line-of-sight (LoS) fluctuations in the $\rm DM$} measurements introduced by the ionization of IGM during EoR. We find that the ratio $\sigma_{\rm DM}/\overline{{\rm DM}}$ depends directly on the ionization bubble sizes as well as the reionization history. On the other hand, we also find that angular variance (coined as structure function) of $\rm DM$} encodes the information about the duration of reionization and the typical bubble sizes as well. We establish the usefulness of variances in $\rm DM$} using toy models of reionization and later verify it with the realistic reionization simulations.
@@ 1 Introduction
According to the current understanding of cosmology, our universe transitioned from being a highly cold-neutral phase in the past to an almost hot-ionized phase at present. This is supposed to be a result of UV radiation from the very first objects that formed in the universe photoionizing the intergalactic medium (IGM) (see e.g., Loeb & Barkana, 2001; Pritchard & Loeb, 2012; Zaroubi, 2013). This window of transition is termed the Epoch of Reionization (EoR). The study of the EoR is crucial for answering several questions regarding the emergence of the first sources, their properties, and impact on the IGM, evolution to present-day structures, the exact timeline of this epoch, etc. Despite our efforts in the last few decades, our understanding of the EoR remains limited (see, e.g., Shaw et al., 2023).
[…]
@@
Probing EoR directly using the redshifted 21 cm signal with the current instruments is also challenging because of several hindrances, such as large ($\sim 10^{4}$ times) foregrounds (e.g., Ali et al., 2008; Bernardi, G. et al., 2009, 2010; Ghosh et al., 2012), thermal noise, radio frequency interference, ionospheric turbulence, and other systematics. While no undisputed detection of the EoR 21 cm signal has been achieved so far, the current data from the radio interferometric experiments have been able to provide a few upper limits on the EoR 21 cm power spectra (e.g., LOFAR: Patil et al. 2017; Mertens et al. 2020, MWA: Barry et al. 2019; Li et al. 2019; Trott et al. 2020, HERA: Abdurashidova et al. 2022, 2023), and the upper limits are improving gradually.
@@
A few earlier works have demonstrated the potential of highly energetic astrophysical events such as gamma-ray bursts (GRBs; e.g., Ciardi & Loeb 2000; Ioka 2003; Inoue 2004; Lidz et al. 2021) and fast radio bursts (FRBs; e.g., Beniamini et al. 2021; Hashimoto et al. 2021; Heimersheim et al. 2022) during EoR as a probe to measure the reionization history. In this work, we focus on the FRBs, which are luminous short-duration (around a few milliseconds) astrophysical radio pulses that have been detected within a frequency band of $0.1-7~{\rm GHz}$ (see Petroff et al., 2022, for a review). Empirical studies on the all-sky event rates of FRBs based on observations find it to be $\sim 10^{3}\,{\rm day}^{-1}$ above the fluence limit of $5$ Jy ms and at a central frequency of $500$ MHz (The CHIME/FRB Collaboration et al., 2021), and the rate is expected to increase significantly at lower fluence thresholds.
@@
FRB signals disperse while traveling through the ionized medium. The amount of dispersion, quantified as Dispersion Measure (DM), directly depends on the free electron content along its path. The DM of a cosmological FRB is expected to have a dominant contribution coming from the electrons in the IGM. During post-EoR ($z\lesssim 5.5$), where the IGM is almost ionized, the IGM DM roughly scales directly with the distances. Therefore, the DM measurements can be turned to infer the redshift distance of the FRB (Zhang, 2018; Kumar & Linder, 2019). Conversely, knowing the redshift of the FRBs accurately can be potentially used to estimate baryonic content of IGM during the post-EoR (e.g., McQuinn, 2013; Macquart et al., 2020; Lee et al., 2022; Khrykin et al., 2024), probe the epoch of second helium reionization (e.g., Caleb et al., 2019; Linder, 2020; Bhattacharya et al., 2021; Lau et al., 2021) and constrain several cosmological parameters (e.g., Deng & Zhang, 2014; Zhou et al., 2014; Yang & Zhang, 2017; Walters et al., 2018; Jaroszynski, 2019; Pol et al., 2019; Wu et al., 2020; Wucknitz et al., 2021; James et al., 2022; Hagstotz et al., 2022). In this work we explore how useful they can be as detailed probes of the EoR.
@@
Despite their enigmatic origin, a recent discovery of a galactic FRB (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020; Bera et al., 2024) clearly associates at least some FRBs with magnetars. Other, less direct, evidence linking FRBs to magnetars comes from the statistical properties of the bursts, from host galaxies and offsets relative to them, and from the energetics and temporal properties of the bursts (e.g., Popov & Postnov, 2010; Falcke, Heino & Rezzolla, Luciano, 2014; Ioka, 2020; Wadiasingh et al., 2020; Zhang, 2022; Beniamini & Kumar, 2023; Chen et al., 2023; Totani & Tsuzuki, 2023). Hence, we can expect a sufficiently large number of FRBs during EoR ($z>6$) which spans a much larger time compared to the lifetime of the massive Population III stars ($\sim~5-30~{\rm Myr}$) that leave behind neutron star (NS) remnants with large angular momentum and strong magnetic fields. There is indirect evidence that supports a relatively large abundance of high-$z$ FRBs (see the introduction of Beniamini et al., 2021, for more details).
@@
With the possibility of detecting high-$z$ FRBs (e.g., Hashimoto et al., 2020; Gupta et al., 2025), one can turn their precise DM measurements to probe the sources and IGM during reionization. Recently, a few theoretical studies (Beniamini et al., 2021; Hashimoto et al., 2021; Heimersheim et al., 2022) have demonstrated the feasibility of using the DM measurements to extract the reionization history and Thomson scattering optical depth ($\tau_{\rm Th}$). Pagano & Fronenberg (2021) and Maity, Barun (2024) have used the mean DM from their synthetic FRB population to constrain the parameters of their reionization simulation. The results in most of these works are based on the assumption of knowing the precise redshift (spectroscopic or empirically) of FRBs. However, detecting the precise spectroscopic redshifts from the host galaxies is a challenging task with the current instruments. Conversely, Beniamini et al. (2021) suggest that the maximum value of DM for bursts spanning the EoR can provide an independent estimate of the Thomson optical depth of the universe without requiring direct redshift information. They have also shown that $\sim 40$ FRBs are sufficient to estimate average electron fraction in four $z$-bins (within $z=6$ to $10$) with $4\%$ accuracy, if their redshifts could be determined within $10\%$ uncertainty. Similar results have been reported in Heimersheim et al. (2022), where they also estimated $\tau_{\rm Th}$ and the mid-redshift of reionization using the average DM. Beniamini et al. (2021) also suggested that the reionization history can be constrained from the determination of the number of FRBs during the EoR per unit DM, i.e., $dN_{\rm FRB}/d{\rm DM}$. Whereas, Hashimoto et al. (2021) have shown that the redshift derivatives of DM have the potential to directly constrain the reionization history.
@@
In this work, we forward model the FRB signal measurements using estimators like globally averaged DM, its redshift derivative, global dispersion, and angular dispersion (along different line-of-sight (LoS)). The primary goal of this work is to study only the DM from IGM, assuming the Milky Way (e.g., Cordes & Lazio, 2002; Yao et al., 2017; Prochaska & Zheng, 2019; Yamasaki & Totani, 2020) and host galaxy (Macquart et al., 2020; Acharya & Beniamini, 2025) contributions are modeled and removed from the observed total DM. However, the models of the host galaxy contributions can differ from reality based on the host galaxy properties, its type, size, inclination, and also the location of the FRB within the galaxy (Mo et al., 2025). These differences can manifest in the IGM contribution as modeling error, which can potentially increase the error bars in the estimated averaged DM. We can expect any modeling error in the host galaxy (at $z>6$) DM would be $\sim 10\,{\rm pc\,cm}^{-3}$, which is roughly $2$ orders of magnitude smaller than the total DM measured for an FRB during EoR. While we agree that modeling of the host DM plays a crucial role for the measurements at low redshifts ($z<1$), it is not so important for DM measurements from reionization. For the first time, we demonstrate that fluctuations along different LoS during EoR potentially bias the mean DM estimates compared to those derived using the mean ionization fraction of the IGM. We also demonstrate that $d{\rm DM}/dz$ and the scatter in DM along different LoS (as a function of $z$) have the potential to discern different reionization histories and morphologies of the IGM. We explore a novel aspect of angular dispersion in DM (defined as structure function in §2.2) at different redshifts. This structure function encodes information regarding the typical bubble sizes in the IGM as it probes the angular fluctuations. We validate our estimators on simple toy models of EoR light-cone (LC) signals. Later, we also apply our estimators to more realistic LCs obtained from simulations. We also perform a comparative study between different reionization histories. For this study, we primarily assume a scenario where the redshifts of the FRBs are known to within $10\%$ uncertainty. Later, we ignore the redshift information and compute the marginalized average DM over the EoR window.
@@
This manuscript is arranged as follows. We define DM and structure function estimators in §2. In §3, we validate our estimators with the toy ionization field model. This is followed by §4 where we briefly describe the details of the actual EoR simulations and present the corresponding results under its subsection, where we discuss the impact of post-EoR IGM on the DM estimates and present results. Finally, we summarize and conclude this exercise in §5. The cosmological simulation here uses the cosmological parameter values $\Omega_{\rm m}=0.27$, $\Omega_{\Lambda}=0.73$, $\Omega_{\rm b}=0.044$ and $h=0.7$ adapted from Hinshaw et al. (2013).
@@ 2 Methodology
We revisit the basic theory of the IGM DM in the context of FRBs and define DM estimators that will be used in this work.
@@ 2.1 Dispersion Measure
The multifrequency FRB pulses disperse while traveling through an ionized medium due to their interaction with the free electrons along the way. The time delay in the arrival of the signal at frequency $\nu$, $\Delta t\propto\nu^{-2}{\rm DM}$; where DM is the line integral of the free electron density. The total observed time delay/DM will have contributions from the host galaxy ($\rm{DM}_{host}$), the Milky Way Galaxy including the halo ($\rm{DM}_{MW}$), and the IGM ($\rm{DM}_{IGM}$}). For $\rm{DM}_{MW}$ we have reasonably good Galactic maps, and this can be largely removed from the data. Further, $\rm{DM}_{host}$ is reduced by a factor of $(1+z)^{-1}$ in the observer frame and so is suppressed when considering high-$z$ events, whereas ${\rm DM}_{\rm IGM}$ increases with $z$. Hence, in this work we only focus on studying $\rm{DM}_{IGM}$} and ignore any further discussion on other DM components, unless stated otherwise.
@@
${\rm{DM}_{IGM}}(\bm{\theta},z)=c\int\limits_{0}^{z}dz^{\prime}\frac{n_{e}(\bm{\theta},z^{\prime})}{(1+z^{\prime})^{2}H(z^{\prime})}~,$ (1)
where $c$ is speed of light in vacuum and $n_{e}(\bm{\theta},z)$ is the proper number density of free electrons. $H(z)=H_{0}\sqrt{\Omega_{m0}(1+z)^{3}+\Omega_{\Lambda 0}}$ denotes the Hubble parameter with $H_{0}$, $\Omega_{m0}$ and $\Omega_{\Lambda 0}$, respectively, being the present-day values of the Hubble constant, matter density parameter, and dark energy parameter. Baryons being the primary source of free electrons during and after EoR, we can write $n_{e}$ in terms of the baryon density parameter $\Omega_{\rm b0}$ and recast eq. (1) as
$\begin{split}{\rm{DM}_{IGM}}(\bm{\theta},z)=\frac{13}{16}&\frac{3cH_{0}\Omega_{\rm b0}}{8\pi Gm_{\rm H}}\times\\ &\int\limits_{0}^{z}dz^{\prime}\frac{(1+z^{\prime})\Delta(\bm{\theta},z^{\prime})x_{\rm{H}\,\textsc{ii}}(\bm{\theta},z^{\prime})}{\sqrt{\Omega_{m0}(1+z^{\prime})^{3}+\Omega_{\Lambda 0}}}~,\end{split}$ (2)
where $G$ is the gravitational constant, $m_{\rm H}$ is the mass of hydrogen atom, $\Delta$ is the matter overdensity and $x_{\rm{H}\,\textsc{ii}}$ denotes the ionization fraction of the IGM. Here, $\Delta$ includes the effects of evolution of the underlying matter density in the IGM, whereas $x_{\rm{H}\,\textsc{ii}}$ is controlled by the photon field responsible for ionizing the IGM. On large scales, $x_{\rm{H}\,\textsc{ii}}=0$ before the reionization starts, and it eventually approaches unity toward the end of EoR when the IGM is almost completely ionized. We obtain both $\Delta$ and $x_{\rm{H}\,\textsc{ii}}$ from our simulations, which we discuss later in §4.1. The treatment of eq. (2) assumes that hydrogen and helium constitute almost the entire baryons in the universe, the helium being $25\%$ of it by mass. It is widely accepted that the photons responsible for the first ionization of He i ($24.6$ eV) are also produced in ample amounts to ionize it concurrently with H i (e.g., Eide et al. 2020). However, the exact ionization histories of hydrogen and helium may differ a little depending on the ionizing source models used. For simplicity, our model of IGM also assumes that the ionization of He i to He ii occurs concurrently with H i reionization.
The LoS path that the FRB signal traverses during post-reionization is effectively very large ($\sim 6000~{\rm Mpc}$), adding a larger contribution to total $\rm{DM}_{IGM}$}. Post-EoR contribution acts as a nuisance since we are only interested in the impact of the reionization on the DM measurements. Hence, we restrict the lower limit of the integral (eq. 2) to the end of reionization $z_{\rm end}$ and define $\rm{DM}_{EoR}$}:
$\begin{split}{\rm{DM}_{EoR}}(\bm{\theta},z)=\frac{13}{16}&\frac{3cH_{0}\Omega_{\rm b0}}{8\pi Gm_{\rm H}}\times\\ &\int\limits_{z_{\rm end}}^{z}dz^{\prime}\frac{(1+z^{\prime})\Delta(\bm{\theta},z^{\prime})x_{\rm{H}\,\textsc{ii}}(\bm{\theta},z^{\prime})}{\sqrt{\Omega_{m0}(1+z^{\prime})^{3}+\Omega_{\Lambda 0}}}~,\end{split}$ (3)